OpenStudy (anonymous):

The point of concurrency of the angle bisectors of a triangle is called the _____ centroid. circumcenter. orthocenter. incenter.

OpenStudy (anonymous):

The length of mc037-2.jpg is shown. GF = 2x + 2 and FE = 4x – 8. What is the value of x? 5 4 12 2

OpenStudy (anonymous):

The concurrency one should be centroid

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem. Glad I could help.

OpenStudy (anonymous):

The length of mc037-2.jpg is shown. GF = 2x + 2 and FE = 4x – 8. What is the value of x? 5 4 12 2

OpenStudy (anonymous):

do you know this problem ?

OpenStudy (anonymous):

There is no picture for it

OpenStudy (anonymous):

this is the pic

OpenStudy (anonymous):

Not sure if I am correct in this, but it looks like the two lines that you told me about intersect. Because they intersect, we can say that at some point on both lines, the equations equal each other.

OpenStudy (anonymous):

it that 5 ??

OpenStudy (anonymous):

Using the method I just explained....yes I got that answer.

OpenStudy (anonymous):

i try that i got 5 is the answer

OpenStudy (anonymous):

Which is what I got

OpenStudy (anonymous):

i mean i got the answer is 5

OpenStudy (anonymous):

what did you get the answer ??

OpenStudy (anonymous):

Like I said, I believe that the answer using my method is 5

OpenStudy (anonymous):

oh ok so we got the same one thank u for ur help

OpenStudy (anonymous):

No problem, glad I could help.

OpenStudy (anonymous):

did you know this one

OpenStudy (anonymous):

Points B, D, and F are midpoints of the sides of mc004-2.jpg. EC = 44 and BF = 2x - 6. Find the value of x. 25 50 14 22

OpenStudy (anonymous):

Same picture or different?

OpenStudy (anonymous):

different one

OpenStudy (anonymous):

Ah interesting...this one.

OpenStudy (anonymous):

14 right

OpenStudy (anonymous):

I believe so